Circular No. 6676

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 K1 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6669), his discovery of a comet in images obtained with the C3 coronagraph aboard the SOHO mission, as the comet moved in toward the sun from 14.6 to 6.4 R_{⊙} over a 23-hr interval, after which a 5-hr span of images was obtained with the C2 coronagraph to a distance of 4.0 R_{⊙}. The following approximate discovery position is extracted from the more complete set of sohocentric J2000.0 positions on MPEC 1997-L02, these having been reduced by G. V. Williams from measurements by St. Cyr and D. A. Biesecker of apparent solar distance and position angle:

1997 UT  α_{2000}  δ_{2000}
May 31.221  4^{h}36.^{m}3  +18^{o}08′

Orbital computations (also on MPEC 1997-L02) by the undersigned confirm St. Cyr’s suggestion that this is another Kreutz sungrazer, with T = 1997 June 1.60 UT (despite some slight inconsistency again between the C3 and C2 data: see IAUC 6650).

COMET 104P/KOWAL 2 A 900-s co-added exposure by C. W. Hergenrother with the 1.2-m

reflector at the Whipple Observatory in mid-May shows a coma 8′′ across and no tail. Red magnitudes: May 14.48 UT, 19.0; 15.42, 19.2.

GRB 970508 M. R. Metzger and J. G. Cohen, California Institute of Technology

(CIT); F. H. Chaffee, W. M. Keck Observatory; and R. D. Blandford, CIT, report: “Images of Bond’s optical variable (IAUC 6654, 6655), associated with GRB 970508 (IAUC 6649), were obtained with the Keck II 10-m telescope (+ LRIS) on June 5.26 UT. The two 5-min R-band exposures give a mean magnitude for the variable of R = 23.4 ± 0.2, tied to the photometric system reported on IAUC 6658. A 30-min spectrum was also obtained (on June 5.29); this reveals an emission line at 684.0 nm that we identify with [O II] 372.8-nm emission at z = 0.835 ± 0.001. This is the same redshift as the strong metal-line absorption system (IAUC 6655), and the spectrum is consistent with having a constant line flux from May 11, but with a considerably weaker continuum. The line flux therefore may be due to nebular emission from a coincident host galaxy, with relatively faint continuum (IAUC 6674).”

1997 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden